. COLLEGE SAVINGS Shannon is choosing between two different savings accounts to
keep her college fund in. The first account compounds interest semiannually at a rate of
11.0%. The second account compounds interest continuously at a rate of 10.8%. If
Shannon plans to keep her money in the account for 5 years, which account should she
choose? Explain.

Found 2 solutions by Boreal, MathTherapy:
Answer by Boreal(15235)

the formula for the first is P=Po(1+(.11/2)^10, the 10 being the number of times compounded.
P=1.708Po
The second is P=Poe^rt, or Poe^(.108*5)or Poe^0.54=1.716 Po
Continuously compounding maximizes return, all other things being equal, and in this instance is better even with a slightly less rate of return.

Answer by MathTherapy(10551)


. COLLEGE SAVINGS Shannon is choosing between two different savings accounts to
keep her college fund in. The first account compounds interest semiannually at a rate of
11.0%. The second account compounds interest continuously at a rate of 10.8%. If
Shannon plans to keep her money in the account for 5 years, which account should she
choose? Explain.
YIELD, based on semi-annual compounding: 
YIELD, based on continuous compounding:
Can you now decide and explain your choice? You need to, because I will not!!